Tanya has a triangular garden bed in her back yard. The measure of the largest angle is 30 degrees less than twice the measure of the smallest. And the measure of the middle angle is 10 degrees more then the smallest. Find the measure of the three angles Start your work by defining the unknown quantities in terms of a variable
Answer by JulietG(1812)

We know that a triangle has 180 degrees.
Angle A + Angle B + Angle C = 180
Let A be the smallest and C be the largest.
C = 2A - 30 (The measure of the largest angle is 30 degrees less than twice the measure of the smallest.)
B = A + 10 (And the measure of the middle angle is 10 degrees more then the smallest.)
Substitute what you know into the first equation.
A + B + C = 180
A + (A+10) + (2A-30) = 180
4A-20 = 180
4A = 200
A = 50
If A = 50, then B = 50+10, and C =2(50)-30
50 + 60 + 70 = 180